OpenStudy (anonymous):

translate each problem into an equation. (1) State College has 620 students. (2) There are 20 more women than men. How many women are there?

OpenStudy (anonymous):

x+y=620

OpenStudy (anonymous):

y =women x=man y=x+20

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do you know what do we have to do next?

OpenStudy (anonymous):

um not quite sure

OpenStudy (anonymous):

x+y=620 y=x+20 We can take the second equation and plug it on the first one. Like this: x+(x+20)=620 How can we continue now?

OpenStudy (anonymous):

distribute your x then solve by adding like terms first

OpenStudy (anonymous):

then 20 - from both sides that leaves 2x=600 now divide

OpenStudy (anonymous):

300?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You got it! Well done!

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

But x is the number of man, right?

OpenStudy (anonymous):

How do we get y now?

OpenStudy (anonymous):

y=x+20 y=300+20 y=320 320 women is the answer.

OpenStudy (anonymous):

Any doubt?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

Awsome! Hope I've helped you! PD: Thanks for the "Best response" :D